# Stochastic Geometry and Dynamical System Analysis of Walker Satellite Constellations

### v2

@@ Abstract
In practice, low Earth orbit (LEO) and medium Earth orbit (MEO) satellite networks consist of multiple orbits which are populated with many satellites. A widely used spatial architecture for LEO or MEO satellites is the Walker constellation, where the longitudes of orbits are equally spaced and the satellites are equally spaced along the orbits. In this paper, we develop a stochastic geometry model for the Walker constellations. This proposed model enables an analysis based on dynamical system theory, which allows one to address essential structural properties such as periodicity and ergodicity. It also enables a stochastic geometry analysis under which we derive the performance of downlink communications of a typical user at a given latitude, as a function of the key constellation parameters.
@@ I-A Motivation and Related Work
LEO and MEO satellite networks are designed to support various applications, including data communications, data sensing and harvesting, and Internet routing. Geometrically, LEO and MEO satellites are located on specific orbits, which determine not only their relative positions but also their motions. Practical satellite networks are primarily designed to maximize coverage, leading to two characteristic geometric features: (i) orbit longitudes are regularly distributed over the equator, and (ii) satellites are regularly placed within each orbit [1]. This configuration is known as the Walker constellation. From the perspective of a fixed observer on Earth, such a Walker constellation is often presumed to lead to periodic ephemerides, given the regular distribution of both orbits and satellites. But is it truly the case?
Recently, stochastic geometry has been widely used to model satellite networks by mathematically characterizing the spatial distribution of satellites. For instance, binomial and Poisson point processes have been employed to represent satellites as uniform random points on a sphere [2, 3, 4, 5]. However, these models fail to capture the orbital structure of practical LEO and MEO satellite networks, where satellites are clustered along orbits. To address this, Cox point processes was proposed to jointly model both the distribution of orbits and the satellites positioned along them [6, 7, 8, 9, 10]. Although the Cox point processes account for orbit-based clustering, they typically assume isotropic configurations, similar to binomial or Poisson models. Consequently, essential dynamical system properties of non-isotropic Walker constellations—such as periodicity and ergodicity—cannot be rigorously studied within the existing stochastic geometry frameworks.
To examine the dynamic of Walker constellations, we develop dedicated stochastic geometry models for them and then analyze their statistical properties. To the best of the authors’ knowledge, this is the first paper that not only formulates a stochastic geometry model for the well-known Walker constellations but also conducts an analysis of its dynamics.
@@ I-B Theoretical Contributions
Using this framework, we derive essential statistical metrics crucial to the performance of the satellite communications. First, we show that the proposed model has a distribution which is invariant with Earth-rotation. Then, we identify the exact analytic condition under which the proposed model is periodic or ergodic. The ergodicity means that the expectation of the snap shots of network performance of the typical user at a given latitude is equal to the empirical average of network performance of any user at the same latitude over a very long time. Then, assuming the typical user associated with its nearest satellite, we derive the distribution of the distance from the typical user to the nearest satellite as a function of orbit inclination, the number of orbits, the number of satellites, and the user latitude, namely $\phi$, $N_{o}$, $N_{s}$, and $l_{u}$, respectively. Moreover, in contrasts to existing stochastic geometry models, the proposed model inherently exhibits an upper-bounded association distance, which allude to the lower bound for the communications from satellites to users on Earth. Assuming a simple antenna gain from satellites, we derive formulas for the Laplace transform of the total interference power and its expectation, respectively. They are crucial for understanding the amount of interference seen by a typical user at any latitude experiences and can therefore be used to design interference mitigation techniques. Finally, we evaluate the coverage probability, namely the signal to interference-plus-noise ratio distribution of the link from the association satellite to the typical user at a given latitude.
@@ II-A Spatial Model and Signal Propagation
In this paper, the three-dimensional space is equipped with the orthonormal basis $(e_{x},e_{y},e_{z})$ at the center of Earth. The reference space is rotating with the Earth’s natural spin, so that a fixed location on Earth has constant coordinates. The $xy$-plane is the equatorial plane, the $x$-axis is the longitude reference axis and the $z$-axis is the North Pole axis.
@@
Let $N_{o}$ be the number of orbits and $N_{s}$ be the number of satellites on each orbit. We assume that the longitudes of the orbits are regularly distributed, namely $N_{o}$ ascending points are distributed regularly on the $[0,2\pi]$ interval. In addition, we assume that at time zero, a uniformly distributed random offset $\bar{\theta}\sim\text{Uniform}(0,2\pi/N_{o})$ is added to the ascending points. Consequently, at time zero, the longitude of the $i$-th orbit ascending point is given by
@@
for $i=1,\ldots,N_{o}$. Note that for all $i$, the same offset value $\bar{\theta}$ is added. This random shift of periodic points leads to longitudinal rotation invariance, or equivalently, invariance of orbit distribution w.r.t. the direction of the Earth’s spin. The inclinations of orbits are assumed to be all equal to $\phi$ and the orbit process at time 0 is denoted by $\operatorname{\mathcal{O}}=\bigcup_{i=1}^{N_{o}}l_{i}=\bigcup_{i=1}^{N_{o}}l(\theta_{i},\phi)$ where $l_{i}(\theta_{i},\phi)$ is the $i$-th orbit, with longitude angle $\theta_{i}$ and inclination angle $\phi.$
Conditionally on these orbits, satellites are distributed regularly on those orbits at time 0, namely the phases of the $N_{s}$ satellites on each orbit are periodically distributed over $2\pi.$ Let $\bar{\omega}\sim\text{Uniform}(0,2\pi/N_{s})$ be an independent random variable. The phase of the $j$-th satellite on the $i$-th orbit at time zero is
@@
for $j=1,\ldots,N_{s}$. In Eq. (2), the same random variable $\bar{\omega}$ is added to the phases of all satellites. Collectively, the satellite point process at time zero is denoted by
@@
where $\psi_{i}$ is the satellite point process on the $i$-th orbit $l_{i}(\theta_{i},\phi)$ and $\delta_{X}$ is the Dirac measure at $X$, and $X_{i,j}$ is the $j$-th satellite of phase angle $\omega_{j}$ on the orbit $l_{i}(\theta_{i},\phi)$.
Then, for $i=1,...,N_{o}$, $j=1,...,N_{s},$ the $(x,y,z)$-coordinates of the satellite at $X_{i,j}$ at time zero are given by
$x$ $=r\sqrt{\cos^{2}(\omega_{j})+\sin^{2}(\omega_{j})\cos^{2}(\phi)}\cos(\hat{\theta}+\theta_{i}),$ (4)
$y$ $=r\sqrt{\cos^{2}(\omega_{j})+\sin^{2}(\omega_{j})\cos^{2}(\phi)}\sin(\hat{\theta}+\theta_{i}),$ (5)
@@
respectively, where $\hat{\theta}=\arctan(\tan(\omega_{j})\cos(\phi))$. The proposed model is a stochastic geometry formulation of the Walker-delta constellation [11], especially when the relative phases between satellites are set to be zero, which can be denoted by Walker $(\phi(\text{rad}):N_{s}\times N_{o},N_{o},0\degree).$
Fig. 1 shows the proposed stochastic geometry model with $N_{o}=20$, $N_{s}=20,$ $\phi=33\degree,$ and $r=7000$ km. The proposed model for the satellite network contrasts with the existing stochastic geometry models based on a binomial point process [2, 4] or a Cox point process [6, 7, 8, 9, 10], where both models are invariant w.r.t. all rotations. In those models, the network users anywhere on Earth see the same satellite distribution, resulting in an average performance metric which is the same for all users on Earth. In contrast, in the proposed model, users at the same latitude observe the same satellite distribution, whereas users at different latitudes experience different satellite distributions. Consequently, network performance metrics vary depending on the user’s latitude.
@@
For performance analysis, all satellites are assumed to use the same wireless resource and network users on Earth are associated with their nearest satellites. Similar to [12, 6.1.3], we assume the received signal power is given by a power-law path loss, multiplied by a random fading and by the antenna gain of the corresponding propagation channel. For the channel of distance $d>1$ (meter), the received signal power is $pG(d)Hd^{-\alpha}$ with $p$ the average isotropic received signal power at $d=1$, $G(d)$ the aggregate antenna gain, and $H$ the random fading over the channel. For $G(d),$ we assume that satellite antennas provide additional gains of $g_{t}$ and the receive antenna provides an isotropic gain of $g_{r}$. Inspired by the practical guideline of [12], we assume the antenna gain over the channel of length $d$ is given by $G(d)=g_{t}g_{r}$ if $d\leq d_{g};$ and $G(d)=g_{r}$ if $d>d_{g}$ where $d_{g}$ is the cutoff distance [^1] This follows from [12] where ground users within a certain angle have an antenna gain $g_{t}$. . The cumulative CDF (CCDF) of $H$ is denoted by $\bar{F}_{H}(x)$ and we assume $\operatorname{\mathbf{E}}[H]=1$.
In this paper, we first examine the dynamical aspect of the proposed stochastic geometry network model. We first derive the distance distribution from the typical user to its nearest satellite, the Laplace transform of signal-plus-interference (SPI) power of the typical user, and then the coverage probability of the typical user, defined as the probability that the signal to interference-plus-noise ratio (SINR) of the typical user is greater than a given threshold $\tau.$
@@ III Dynamical System Analysis
Let $S$ be a measurable space. A dynamical system $R$ with state space $S$ is a deterministic measurable map $R$ from $S\times\mathbb{R}^{+}$ to $S$. The state of the dynamical system at time $t$ is $R(x,t)$, with $x$ the initial state, namely $R(x,0)=x$. It is said to admit an invariant measure if there exists a probability measure $\cal P$ on $S$ such that if $x$ is chosen at random with the distribution $\cal P$, then $R(x,t)$ has $\cal P$ for distribution for all $t\geq 0$. It is said to be minimal if, for all $x\in S$, the trajectory $R(x,t)$ is dense in $S$. Let $h:S\to\mathbb{R}$ be a measurable function. The time average of $h(R)$ associated with the initial condition $x$ is defined as $\lim_{T\to\infty}\frac{1}{T}\int_{0}^{T}h(R(x,t))\mathrm{d}t$, when the limit exists. If $\cal P$ is an invariant probability measure for $R$, then, for all measurable functions $h:S\to\mathbb{R}$ such that $\int|h(s)|{\cal P}(\mathrm{d}s)<\infty$, the time average of $h(R)$ exists for $\cal P$ almost-all $x$ (this is Birkhoff’s theorem [13]). The dynamical system $R$ is said to be ergodic w.r.t. $\cal P$ if, for all such $h$, and for $\cal P$ almost-all $x$,
$\lim_{T\to\infty}\frac{1}{T}\int_{0}^{T}h(R(x,t))\mathrm{d}t=\int h(s){\cal P}(\mathrm{d}s),$ (7)
i.e., the time average of $h(R)$ is the mean of $h(R)$ w.r.t. $\cal P$.
@@
Consider the dynamical system $R_{\alpha}(x,t)$ on $[0,1]$, where $R_{\alpha}(x,t)=x+\alpha t\mod 1$. Here, $x$ is the initial condition, $\alpha$ is the velocity, and $R_{\alpha}(x,t)$ is the state at time $t$. Then, the uniform measure on $[0,1]$, $\cal P$, is an invariant probability measure of $R_{\alpha}$. If $\alpha$ is rational, the dynamic is periodic and is hence neither minimal nor ergodic w.r.t. $\cal P$. If $\alpha$ is irrational, the dynamical system is minimal and ergodic w.r.t. $\cal P$.
In the proposed satellite model, let $\bar{\theta}(t)$ denote the longitude of the first orbit at time $t$ and $\bar{\omega}(t)$ the angle (mesured in the orbital plane) of the first satellite on this orbit at time $t$. Let $\bar{\theta}$ and $\bar{\omega}$ denote the initial condition of these state variables. Then,
$\bar{\theta}(t)=\bar{\theta}-\bar{v}_{{\theta}}t\mod\frac{2\pi}{N_{0}}\text{ and }\bar{\omega}(t)=\bar{\omega}+\bar{v}_{{\omega}}t\mod\frac{2\pi}{N_{s}},$
where $\bar{v}_{\theta}$ and $\bar{v}_{\omega}$ denote the angular speed of Earth-spin and the angular speed of satellites’ rotation, respectively. Our key dynamical system is $R(x,t)=(\bar{\theta}(t),\bar{\omega}(t))$, with state space $S=[0,\frac{2\pi}{N_{o}})\times[0,\frac{2\pi}{N_{s}})$ and with initial condition $x=(\bar{\theta},\bar{\omega})$.
We now describe the main ideas developed in the forthcoming subsections. Let $\Psi(x,t)$ denote the Walker satellite point process at time $t$, namely the point process $\Psi$ in Eq. (3), with $\bar{\theta}$ replaced by $\bar{\theta}(t)$ and $\bar{\omega}$ replaced by $\bar{\omega}(t)$, respectively. Note that $\Psi(x,t)$ is a deterministic measurable functional $\Xi$ of the key dynamical system $R(x,t)$. Therefore, if $R$ has an invariant probability measure $\cal P$ on $S$, then $\Psi$ has an invariant probability measure as well, which is the push-forward of $\cal P$ by $\Xi$, denoted by ${\cal P}_{\Psi}$, which is a probability measure on the space $\mathcal{N}_{r}$ of counting measures on the sphere of radius $r$. Note that ${\cal P}_{\Psi}$ is just the distribution of $\Psi$ at time zero when $(\bar{\theta},\bar{\omega})$ has the distribution $\cal P$. In addition, if $R$ is ergodic w.r.t. $\cal P$, so is the dynamical system $\Psi$ w.r.t. ${\cal P}_{\Psi}$. Since this point process is described in the referential space that rotates with Earth, in this referential, a static Earth user is a point with static coordinates. Hence, $\Psi(x,t)$ provides a complete description of the ephemerides of the Walker constellation.

### v3

@@ Abstract
In practice, low Earth orbit (LEO) and medium Earth orbit (MEO) satellite networks consist of multiple orbits which are populated with many satellites. A widely used spatial architecture for LEO or MEO satellites is the Walker constellation, where the longitudes of orbits are evenly spaced and the satellites are equally spaced along the orbits. In this paper, we develop a stochastic geometry model for the Walker constellations. This proposed model enables an analysis based on dynamical system theory, which allows one to address essential structural properties such as periodicity and ergodicity. It also enables a stochastic geometry analysis under which we derive the performance of downlink communications of a typical user at a given latitude, as a function of the key constellation parameters.
@@ I-A Motivation and Related Work
LEO and MEO satellite networks are designed to support various applications, including data communications, data sensing and harvesting, and Internet routing. Geometrically, LEO and MEO satellites are located on specific orbits, which determine not only their relative positions but also their motions. Practical satellite networks are primarily designed to maximize coverage based on two characteristic geometric features: (i) orbit longitudes are distributed evenly over the equator, and (ii) satellites are placed evenly within each orbit [1]. This configuration is known as the Walker constellation. From the perspective of a fixed observer on Earth, such a Walker constellation is often presumed to lead to periodic ephemerides, given the regular distribution of both orbits and satellites. But is it truly the case?
Recently, stochastic geometry has been widely used to model satellite networks by mathematically characterizing the spatial distribution of satellites. For instance, binomial and Poisson point processes have been employed to represent satellites as uniform random points on a sphere [2, 3]. However, these models fail to capture the orbital structure of practical LEO and MEO satellite networks, where satellites are clustered along orbits. To address this, Cox point processes were proposed to jointly model both the distribution of orbits and the satellites positioned along them [4, 5, 6, 7, 8]. Although Cox point processes account for orbit-based clustering, they typically assume isotropic configurations, similar to binomial or Poisson models. Consequently, essential dynamical system properties of non-isotropic Walker constellations—such as periodicity and ergodicity—cannot be rigorously studied within the existing stochastic geometry frameworks. To examine the dynamics of Walker constellations, we develop dedicated stochastic geometry models for them and then analyze their statistical properties.
@@ I-B Theoretical Contributions
Using this framework, we derive essential statistical metrics crucial to the performance of the satellite communications. First, we show that the proposed model has a distribution which is invariant with time. Then, we identify the exact analytic condition under which the proposed model is periodic or ergodic. The ergodicity means that the expectation of a snap shots of network performance experienced by the typical user at a given latitude is equal to the empirical average of network performance of any user at the same latitude over a very long time. Then, assuming the typical user associates with its nearest satellite, we derive the distribution of the distance from the typical user to its serving satellite as a function of orbit inclination, the number of orbits, the number of satellites, and the user latitude, namely $\phi$, $N_{o}$, $N_{s}$, and $l_{u}$, respectively. Moreover, in contrasts to existing stochastic geometry models, the proposed model inherently exhibits an upper-bounded association distance, which, under appropriate conditions, lead to a lower bound on the rates of communications from satellites to users on Earth. Assuming a simple antenna gain model for satellites, we derive formulas for the Laplace transform of the total interference power and its expectation, respectively. They are crucial for understanding the amount of interference seen by a typical user at any latitude and therefore can be used to design interference mitigation techniques. Finally, we evaluate the coverage probability, namely the signal to interference-plus-noise ratio (SINR) distribution of the link from the association satellite to the typical user at a given latitude.
@@ II-A Spatial Model and Signal Propagation
In this paper, the three-dimensional space is equipped with the orthonormal basis $(e_{x},e_{y},e_{z})$ with the origin at the center of Earth. The reference space is rotating with the Earth’s natural spin, so that a fixed location on Earth has constant coordinates. The $xy$-plane is the equatorial plane, the $x$-axis is the longitude reference axis and the $z$-axis is the North Pole axis.
@@
Let $N_{o}$ be the number of orbits and $N_{s}$ be the number of satellites on each orbit. We assume that the longitudes of the orbits are evenly distributed, namely those $N_{o}$ ascending points are distributed evenly on the $[0,2\pi]$ interval. In addition, we assume that, at time zero, a uniformly distributed random offset $\bar{\theta}\sim\text{Uniform}(0,2\pi/N_{o})$ is added to the ascending points. Consequently, at time zero, the longitude of the $i$-th orbit ascending point is given by
@@
for $i=1,\ldots,N_{o}$. Note that for all $i$, the same offset $\bar{\theta}$ is added. This random shift of periodic points leads to longitudinal rotation invariance, or equivalently, invariance of orbit distribution w.r.t. the Earth’s spin. The inclinations of orbits are assumed to be all equal to $\phi$ and the orbit process at time 0 is denoted by $\operatorname{\mathcal{O}}=\bigcup_{i=1}^{N_{o}}l_{i}=\bigcup_{i=1}^{N_{o}}l(\theta_{i},\phi)$ where $l_{i}(\theta_{i},\phi)$ is the $i$-th orbit, with longitude angle $\theta_{i}$ and inclination angle $\phi.$
Conditionally on these orbits, satellites are distributed regularly on each orbit at time 0, namely the phases of the $N_{s}$ satellites on each orbit are evenly distributed over $[0,2\pi].$ Let $\bar{\omega}\sim\text{Uniform}(0,2\pi/N_{s})$ be an independent random variable. The phase of the $j$-th satellite on the $i$-th orbit at time zero is
@@
for $j=1,\ldots,N_{s}$. In Eq. (2), the same random variable $\bar{\omega}$ is added to the phases of all satellites. Collectively, the satellite point process at time zero is given by
@@
where $\psi_{i}$ is the satellite point process on the $i$-th orbit $l_{i}(\theta_{i},\phi)$ and $\delta_{X}$ is the Dirac measure at $X$, and $X_{i,j}$ is the $j$-th satellite with its phase $\omega_{j}$ on the orbit $l_{i}(\theta_{i},\phi)$.
[…]
$x$ $=r\sqrt{\cos^{2}(\omega_{j})+\sin^{2}(\omega_{j})\cos^{2}(\phi)}\cos(\hat{\theta}_{j}+\theta_{i}),$ (4)
$y$ $=r\sqrt{\cos^{2}(\omega_{j})+\sin^{2}(\omega_{j})\cos^{2}(\phi)}\sin(\hat{\theta}_{j}+\theta_{i}),$ (5)
@@
Fig. 1 depicts the proposed stochastic geometry model with $N_{o}=20$, $N_{s}=20,$ $\phi=33\degree,$ and $r=7000$ km. The proposed model for the satellite network contrasts with the existing stochastic geometry models based on binomial point processes [2] or an isotropic Cox point process [4, 5, 6, 7, 8], where both models have distributions invariant w.r.t. all rotations. In those models, the network users anywhere on Earth see the same satellite distribution, resulting in an average performance metric which is the same for all users on Earth, regardless of the users’ latitudes. In contrast, in the proposed model, users at different latitudes experience different satellite distributions. Consequently, network performance metrics vary depending on the user’s latitude. This aspect will be explained further.
[…]
For performance analysis, all satellites are assumed to use the same wireless resource and network users on Earth are associated with their nearest satellites. As in [11, 6.1.3], we assume the received signal power is given by a power-law path loss, multiplied by a random fading and by the antenna gain of the corresponding propagation channel. For the channel of distance $d>1$ (meter), the received signal power is $pG(d)Hd^{-\alpha}$ with $p$ the average isotropic received signal power at $d=1$, $G(d)$ the aggregate antenna gain, and $H$ the random fading over the channel. For $G(d),$ we assume that satellite antennas provide additional gains of $g_{t}$ and the receive antenna provides an isotropic gain of $g_{r}$. Inspired by the practical guideline in [11], we assume the antenna gain over the channel of length $d$ is given by $G(d)=g_{t}g_{r}$ if $d\leq d_{g};$ and $G(d)=g_{r}$ if $d>d_{g}$ where $d_{g}$ is the cutoff distance [^1] This follows from [11] where ground users within a certain angle have an antenna gain $g_{t}$. . The cumulative CDF (CCDF) of $H$ is denoted by $\bar{F}_{H}(x)$ and we assume $\operatorname{\mathbf{E}}[H]=1$.
In this paper, we first examine the dynamical aspect of the proposed stochastic geometry network model. We then derive the distance distribution from the typical user to its nearest satellite, the Laplace transform of signal-plus-interference power of the typical user, and finally the coverage probability of the typical user, defined as the probability that the SINR of the typical user is greater than a given threshold $\tau.$
@@ III Dynamical System Analysis
Let $S$ be a measurable space. A dynamical system on $S$ is a measurable flow $\{R_{t}\}_{t\in\mathbb{R}^{+}}$ on $S$, namely a family of deterministic maps $R_{t}$ from $S$ to itself, such that (i) the map $(t,x)\to R_{t}(x)$ is measurable from +×$\mathbb{R}^{+}\times S$ to $S$, and (ii) $R_{t}\circ R_{s}=R_{t+s}$, for all $s,t\in\mathbb{R}^{+}$. The state of the dynamical system at time $t$ is $R_{t}(x)$, with $x$ the initial state, namely $R_{0}(x)=x$.
Moreover, it is said to admit an invariant measure if there exists a probability measure $\mathcal{P}$ on $S$ such that if $x$ is chosen at random with the distribution $\mathcal{P}$, then $R_{t}(x)$ has $\mathcal{P}$ for distribution for all $t\geq 0$.
In case $S$ is a metric space equipped with its Borel $\sigma$-algebra and the map $(t,x)\to R_{t}(x)$ is continuous, the flow is said to be minimal if, the trajectory $R_{t}(x)$ is dense in $S$ for all $x\in S$.
Let $\{R_{t}\}$ be a measurable flow and let $h:S\to\mathbb{R}$ be a measurable function. The time average of $h(R)$ associated with the initial condition $x$ is defined as $\lim_{T\to\infty}\frac{1}{T}\int_{0}^{T}h(R_{t}(x))\mathrm{d}t$, when the limit exists. If $\mathcal{P}$ is an invariant probability measure for $R$, then, for all measurable functions $h:S\to\mathbb{R}$ such that $\int|h(s)|{\mathcal{P}}(\mathrm{d}s)<\infty$, the time average of $h(R)$ exists for $\mathcal{P}$ almost-all $x$ (this is Birkhoff’s theorem [12]). The dynamical system $R$ is said to be ergodic w.r.t. $\mathcal{P}$ if, for all such $h$, and for $\mathcal{P}$ almost-all $x$,
$\lim_{T\to\infty}\frac{1}{T}\int_{0}^{T}h(R_{t}(x))\mathrm{d}t=\int h(s){\mathcal{P}}(\mathrm{d}s),$ (7)
i.e., the long term time average of $h(R)$ is the mean of $h(R)$ w.r.t. the invariant distribution $\mathcal{P}$.
@@
Consider the measurable flow $R^{\alpha}_{t}(x)$ on $S=[0,1]$, where $R_{t}^{\alpha}(x)=x+\alpha t\mod 1$. Here, $x$ is the initial condition, $\alpha$ is the velocity, and $R_{t}^{\alpha}(x)$ is the state at time $t$. Note that $S$ is a compact metric space and that this flow is continuous. The uniform measure on [0,1][0,1][ 0 , 1 ], $\mathcal{P}$, is an invariant probability measure of $R^{\alpha}$. If $\alpha$ is rational, the dynamic is periodic and is hence neither minimal nor ergodic w.r.t. $\mathcal{P}$. If $\alpha$ is irrational, then the dynamical system is minimal and ergodic w.r.t. $\mathcal{P}$.
In the proposed Walker satellite model, let $\bar{\theta}_{t}$ denote the longitude of the first orbit (that with the smallest longitude) at time $t$ and $\bar{\omega}_{t}$ the angle of the first satellite on this orbit (that with the smallest phase) at time $t$. Let $\bar{\theta}$ and $\bar{\omega}$ denote the initial condition of these state variables. Then,
$\bar{\theta}_{t}(\bar{\theta})=\bar{\theta}-\bar{v}_{{\theta}}t\hskip-5.69054pt\mod\frac{2\pi}{N_{0}}\text{ and }\bar{\omega}_{t}(\bar{\omega})=\bar{\omega}+\bar{v}_{{\omega}}t\hskip-5.69054pt\mod\frac{2\pi}{N_{s}},$
where $\bar{v}_{\theta}$ and $\bar{v}_{\omega}$ denote the angular speed of Earth-spin and the angular speed of satellite rotation, respectively. Our key measurable flow is $R_{t}(x)=(\bar{\theta}_{t}(\bar{\theta}),\bar{\omega}_{t}(\bar{\omega}))$, with state space $S=[0,\frac{2\pi}{N_{o}})\times[0,\frac{2\pi}{N_{s}})$ and with initial condition $x=(\bar{\theta},\bar{\omega})$.
We now describe the main ideas developed in the forthcoming subsections. Let $\Psi_{t}$ denote the satellite point process at time $t$, namely the point process $\Psi$ in Eq. (3), with $\bar{\theta}$ replaced by $\bar{\theta}_{t}(\bar{\theta})$ and $\bar{\omega}$ replaced by $\bar{\omega}_{t}(\bar{\omega})$, respectively. Since this point process is described in the referential space that rotates with Earth, in this referential, a static Earth user is a point with static coordinates. Hence, $\Psi_{t}$ provides a complete description of the ephemerides of the Walker constellation. The space $\mathcal{N}_{r}$ of counting measures, namely the sphere of radius $r$ where the satellites move, is a measurable space.
Here are two important observations: (i) $\{\Psi_{t}\}$ is a measurable flow on $\mathcal{N}_{r}$; (ii) for all $t\geq 0$, $\Psi_{t}$ is a deterministic measurable functional $\Xi$ of $R_{t}$. The map $\Xi$ is measurable from $S$ to $\mathcal{N}_{r}$. Therefore, if $\{R_{t}\}$ has an invariant probability measure $\mathcal{P}$ on $S$, then $\{\Psi_{t}\}$ has an invariant probability measure as well, which is the push-forward of $\mathcal{P}$ by $\Xi$ on $\mathcal{N}_{r}$, denoted by ${\mathcal{P}}_{\Psi}$, which is a probability measure on $\mathcal{N}_{r}$. Note that ${\mathcal{P}}_{\Psi}$ is the distribution of $\Psi_{0}$ when $(\bar{\theta},\bar{\omega})$ has the distribution $\mathcal{P}$. In addition, if $\{R_{t}\}$ is ergodic w.r.t. $\mathcal{P}$, then $\{\Psi_{t}\}$ is ergodic w.r.t. ${\mathcal{P}}_{\Psi}$. These last properties can also be understood as the measurable conjugacy of the dynamical systems $\{R_{t}\}$ and $\{\Psi_{t}\}$ (see, e.g., Definition 6.8.2 in [13]).